# Contrastive learning of T cell receptor representations

### v1

@@ III.3.2 Procedure
III.4 Using SVCs to learn PLM features
To train the linear SVCs on top of PLM features, sampled 1000 random background TCRs from the training partition of the unlabelled Tanno et al. dataset. For each PLM-pMHC-split combination, we trained a linear SVC using the PLM embeddings of the reference TCRs as the positives and those of the 1000 background TCRs as the negatives. The same 1000 background TCRs were used across model-pMHC-split combinations to ensure consistency. We accounted for the imbalance between the number of positive and negative samples used during SVC fitting by weighting the penalty contributions accordingly. Finally, we tested SVCs using the same benchmarking classification task as previously described.
III.5 SCEPTR fine-tuning with supervised contrastive learning
III.5.1 Data
For supervised contrastive fine-tuning we took all TCR binders against the six best-sampled pMHC targets from our labelled TCR dataset, and split them into a training, a validation, and a test set such that no study used to construct the training or validation sets contributed any TCRs to the test set (Table SII.
III.5.2 Procedure
@@ Acknowledgements.
The authors thank Ned Wingreen, Chris Watkins, Linda Li, Rudy Yuen, Sankalan Bhattacharyya, and Matthew Cowley for useful discussions. YN and MM were supported by Cancer Research UK studentships under grants BCCG1C8R and A29287, respectively. The work of ATM was supported in parts by funding by the Royal Free Charity.Code Availability
@@
This is a readily usable deployment of SCEPTR and all its variants seen in this publication.
@@
This repository houses code used for designing and training our models.
@@
This repository houses some library code that powers both the sceptr and tcrlm repositories above.
@@ A.3 ProtBert
Appendix B Effects of different position embedding methods

### v2

@@ III.3.2 Procedure
III.4 SCEPTR fine-tuning with supervised contrastive learning
III.4.1 Data
For supervised contrastive fine-tuning we took all TCR binders against the six best-sampled pMHC targets from our labelled TCR dataset, and split them into a training, a validation, and a test set such that no study used to construct the training or validation sets contributed any TCRs to the test set (Table SII).
III.4.2 Procedure
@@ Acknowledgements.
The authors thank Ned Wingreen, Chris Watkins, Trevor Graham, Sergio Quezada, Machel Reid, Linda Li, Rudy Yuen, Sankalan Bhattacharyya, and Matthew Cowley for useful discussions. YN and MM were supported by Cancer Research UK studentships under grants BCCG1C8R and A29287, respectively. The work of ATM was supported in parts by funding by the Royal Free Charity. The authors declare no competing interests.Code Availability
@@
a readily usable deployment of SCEPTR and its variants.
@@
houses code used for designing and training our models.
@@
library code powering the above repositories
@@ A.3 ProtBert
Appendix B Learning features within embedding spaces
The focus of the current work has been to use nearest neighbour prediction using PLM embeddings as the most direct test of data-efficient transfer learning that works with as little as a single reference sequence. If slightly more data is available, another approach is to train supervised predictors atop PLM embeddings. To test how much such training can improve prediction performance, we trained linear support vector classifiers (SVC) on the PLM embeddings provided by different models. In each instance, we trained the classifier to distinguish reference TCRs from 1000 randomly sampled background TCRs. We outline the methodology in more detail below.
We find that the SVC predictors for ProtBert, ESM2 and TCR-BERT all perform better than their nearest neighbour counterparts, but still worse than SCEPTR’s nearest neighbour predictions (Fig. S4). We also trained an SVC atop SCEPTR, which did not lead to further improvement upon the nearest neighbour prediction (Fig. S4). These findings highlight how in the low data regime typical of most pMHCs, misalignment of pre-training to downstream tasks can only be partially remediated by training on reference TCRs.
To train the linear SVCs on top of PLM features, we sampled 1000 random background TCRs from the training partition of the unlabelled Tanno et al. dataset. We employed a similar strategy to our benchmarking in section I.1 to split our dataset of curated specificity-annotated TCRs into a reference set and testing set. For each PLM-pMHC-split combination, we trained a linear SVC using the PLM embeddings of the reference TCRs as the positives and those of the 1000 background TCRs as the negatives. The same 1000 background TCRs were used across model-pMHC-split combinations to ensure consistency. We accounted for the imbalance between the number of positive and negative samples used during SVC fitting by weighting the penalty contributions accordingly. Finally, we tested SVCs using the same benchmarking classification task as previously described.
Appendix C Effects of different position embedding methods